# PRobELM: Plausibility Ranking Evaluation for Language Models

### v1

@@ 2 Related Work
Factual errors in an LLM’s output are also referred to as hallucinations (Xu et al., 2024) and occur when models generate content that is unfaithful to the provided input, established context, or factual correctness. Ji et al. (2023) categorises these into input-conflicting, context-conflicting, and fact-conflicting hallucinations, with the latter category being particularly problematic due to its contradiction with known facts. While all scenarios in the PRobELM benchmark are technically fact-conflicting hallucinations (as they are not yet facts of the world the LLM has seen), the crucial point is that under existing world knowledge, non-factual scenarios are not all equally plausible. Instead, PRobELM probes a model’s ability to make educated guesses about plausible futures to gain insights into the capabilities of LLMs to infer within the limits of their knowledge beyond recalling facts it has already seen.
@@ 4.1 Experimental Setup
We first collect data points from Wikidata that align the timeframe of a given model (outlined in Table 1), following the steps detailed in section 3. For each most plausible scenario, we always add 10 less plausible alternatives, resulting in a total of 11 scenarios for models to rank. To ensure our evaluation is comprehensive yet with the practical constraints of time and computational resources, we sampled 126,000 scenarios (6,000 most plausible scenarios and 120,000 less plausible scenarios) from each timestamp, leading to four distinct timeframe-specific evaluation datasets. This sampling covered the five models under study, with OLMo and LLaMA models sharing the same timestamp. Details of the models and the corresponding timestamps are included in Appendix A. Next, we follow the quality control measurements as detailed in subsection 3.3 to filter our data points. The final number of samples for each timestamp varies, as shown in Table 1.
@@ 5.1 Main Results on PRobELM
The results revealed that while model performance exceeded random chance, overall scores remained low, underscoring the complexity of the plausibility assessment task. For instance, considering a random baseline accuracy of 9.09% when ranking 11 scenarios, most models surpassed this baseline across various sizes and prompts. Notably, both BLOOM models marginally outperformed the random baseline by approximately 3 points, achieving 12.54% and 12.80% on the question-answering prompt. Specifically, the BLOOM-560M model underperformed relative to the random baseline in the text completion prompt with 8.78% accuracy. This performance pattern suggests that language models face significant challenges in differentiating between scenarios of varying plausibility. It also highlights a distinct separation between models’ capabilities in factual accuracy and their ability to infer plausibility. These observations set the stage for our key findings:
@@
While variations in performance across prompts and metrics are observed, they typically correlate strongly. The most significant gap occurred between OLMo-1B’s performance on the text completion prompt (22.11 acc.) and the question answering prompt (42.15 acc.), highlighting some variation in model proficiency across different types of tasks. Despite these variations, the assessment appears to hinge on a consistent underlying mechanism across prompt types. This is further evidenced by the Pearson correlation coefficients between the performances on various prompts, which were notably high: 0.8281 for statement versus text completion, 0.8672 for statement versus question answering, and 0.8976 for text completion versus question answering. Such strong correlations underscore that, despite the prompt-specific performance disparities, the models are evaluated based on a core aspect of their capability of ranking plausibility. Similarly, the Pearson correlation score between MRR and NDCG is 0.9982, indicating consistent plausibility measurements across metrics.
@@ 5.2 Comparison with Other Benchmarks
PRobELM shows distinct model performance compared to other reasoning benchmarks. Models that perform well on tasks emphasising factual accuracy do not always maintain their performance levels on plausibility assessments. In some cases, the performance rankings were inversely related. A notable example of this phenomenon is observed in the performance disparity between PRobELM and TruthfulQA. Specifically, Pythia-14M, which exhibits the best performance in TruthfulQA, positions near the bottom (9th place) of the PRobELM ranking. This variance underscores the distinct nature of plausibility as an evaluation criterion, demanding separate capabilities beyond those required for tasks focused on factual accuracy. The development of PRobELM thus contributes to the broader framework for evaluating language models by offering a different angle to assess their capability to understand and generate plausible content.
Figure 2: Plausibility scores of language models on the time-frame specific datasets. We illustrate one model from each family. Vertical dotted lines represent the training timeframe of each model.
@@ 5.3 Temporal Effect of PRobELM
We further evaluate models against datasets from 2017, 2020, 2021, and 2023 timestamps to investigate how the recency of training data affects the plausibility performance. The outcomes of this analysis are presented in Figure 2 and detailed in Appendix C. Remarkably, the relative rankings of models proved relatively consistent across these varied timeframes, affirming the robustness of our evaluation approach. We observed that, from the 2020 to 2023 timestamps, models trained closer to or after the evaluation dates typically exhibit enhanced performance, underscoring the benefits of leveraging recent data.
@@ Appendix A Models Used in Our Experiments
(Touvron et al., 2023) LLaMA is recognized for its robust performance across a broad range of NLP tasks. Its training data, current as of August 2022, positions it uniquely for evaluation with the PRobELM benchmark, specifically designed to assess the model’s ability to infer plausibility based on its training up to that point. The closest corresponding timestamp we found is from December 2023 to January 2023.
@@ Appendix D More details on PRobELM
P5096
member of the crew of
<subject>is a member of the crew of <object>.
P122
basic form of government
The basic form of government of <subject>is <object>.
P3448
stepparent
<subject>is the stepparent of <object>.
P1479
has contributing factor
<object>is a contributing factor to <subject>.
P61
discoverer or inventor
<subject>is the discoverer or inventor of <object>.
P3320
board member
<subject>serves as a board member of <object>.
P7779
member of military unit
<subject>is a member of the military unit <object>.
P98
editor
<subject>is the editor of <object>.
P1411
nominated for
<subject>was nominated for <object>.
P371
presenter
<subject>is the presenter of <object>.
P1365
replaces
<subject>replaces <object>.
P488
chairperson
<subject>serves as the chairperson of <object>.
P27
country of citizenship
<subject>’s country of citizenship is <object>.
P20
place of death
<subject>’s place of death is <object>.
P1344
participant in
<subject>was a participant in <object>.
P1366
replaced by
<subject>was replaced by <object>.
P1412
languages spoken written or signed
@@
<subject>speaks the following languages: <object>.
P276
location
The location of <subject>is <object>.
P407
language of work or name
The language of <subject>’s work or name is <object>.
P39
position held
<subject>holds the position of <object>.
P1532
country for sport
<subject>represents <object>in sports competitions.
P451
unmarried partner
<subject>is the unmarried partner of <object>.
P54
member of sports team
<subject>is a member of the sports team <object>.
P800
notable work
<subject>’s notable work includes <object>.
P551
residence
<subject>’s residence is in <object>.
P131
located in the administrative territorial entity
<subject>is located in the administrative territorial entity <object>.
P106
occupation
<subject>’s occupation is <object>.
P69
educated at
<subject>was educated at <object>.
P509
cause of death
<subject>’s cause of death was <object>.
P102
member of political party
<subject>is a member of the political party <object>.
P19
place of birth
<subject>’s place of birth is <object>.
P115
home venue
<subject>’s home venue is <object>.
P1001
applies to jurisdiction
<object>applies to the jurisdiction of <subject>.
P840
narrative location
<subject>is set in the narrative location of <object>.
P108
employer
<subject>is employed by <object>.
P57
director
<subject>is the director of <object>.
P2416
sports discipline competed in
<subject>competes in the sports discipline of <object>.
P400
platform
<subject>is available on the platform <object>.
P1433
published in
<subject>was published in <object>.
P1056
product or material produced
<subject>produces <object>as a product or material.
P9071
character type
<subject>is characterized as a <object>type.
P4100
parliamentary group
<subject>is a member of the parliamentary group <object>.
P937
work location
<subject>’s work location is <object>.
P1066
student of
<subject>is a student of <object>.
P1535
used by
<object>is used by <subject>.
P6
head of government
<subject>is the head of government of <object>.
P2283
use
<subject>is used for <object>.
P812
academic major
<subject>’s academic major is <object>.
P1416
affiliation
<subject>is affiliated with <object>.
P2522
victory
<subject>achieved a victory in <object>.
P607
conflict
<subject>was involved in the conflict <object>.
P749
parent organization
<subject>is a part of the parent organization <object>.
P2283
uses
<subject>uses <object>.
P802
student
<subject>is a student at <object>.
P119
place of burial
<subject>’s place of burial is <object>.
P2842
place of marriage
<subject>was married at <object>.
P286
head coach
<subject>is the head coach of <object>.
P2541
operating area
<subject>’s operating area is <object>.
P1441
present in work
<subject>is present in the work <object>.
P2650
interested in
<subject>is interested in <object>.
P1027
conferred by
<object>is conferred by <subject>.
P3300
musical conductor
<subject>is the musical conductor of <object>.
P2715
elected in
<subject>was elected in <object>.
P2937
parliamentary term
<subject>served during the parliamentary term <object>.
P1399
convicted of
<subject>was convicted of <object>.
P1686
for work
<subject>is used for the work <object>.
P1196
manner of death
<subject>’s manner of death was <object>.
P2632
place of detention
<subject>was detained at <object>.
P991
successful candidate
<subject>was the successful candidate in <object>.
P2443
stage reached
<subject>reached the stage <object>.
P6872
has written for
<subject>has written for <object>.
@@
P5096
member of the crew of
Fill in the blank: <subject>is a member of the crew of ____. Answer: <object>
P122
basic form of government
@@
Fill in the blank: The basic form of government of <subject>is ____. Answer: <object>
P3448
stepparent
Fill in the blank: <subject>is the stepparent of ____. Answer: <object>
P1479
has contributing factor
Fill in the blank: <object>is a contributing factor to ____. Answer: <subject>
P61
discoverer or inventor
Fill in the blank: ____ is the discoverer or inventor of <object>. Answer: <subject>
P3320
board member
@@
Fill in the blank: <subject>serves as a board member of ____. Answer: <object>
P7779
member of military unit
@@
Fill in the blank: <subject>is a member of the military unit ____. Answer: <object>
P98
editor
Fill in the blank: ____ is the editor of <object>. Answer: <subject>
P1411
nominated for
Fill in the blank: ____ was nominated for <object>. Answer: <subject>
P371
presenter
Fill in the blank: ____ is the presenter of <object>. Answer: <subject>
P1365
replaces
Fill in the blank: <subject>replaces ____. Answer: <object>
P488
chairperson
Fill in the blank: ____ serves as the chairperson of <object>. Answer: <subject>
P27
country of citizenship
Fill in the blank: <subject>’s country of citizenship is ____. Answer: <object>
P20
place of death
Fill in the blank: <subject>’s place of death is ____. Answer: <object>
P1344
participant in
Fill in the blank: ____ was a participant in <object>. Answer: <subject>
P1366
replaced by
Fill in the blank: <subject>was replaced by ____. Answer: <object>
P1412
languages spoken written or signed
@@
Fill in the blank: <subject>speaks the following languages: ____. Answer: <object>
P276
location
Fill in the blank: The location of <subject>is ____. Answer: <object>
P407
language of work or name
@@
Fill in the blank: The language of <subject>’s work or name is ____. Answer: <object>
P39
position held
@@
Fill in the blank: <subject>holds the position of ____. Answer: <object>
P1532
country for sport
@@
Fill in the blank: ____ represents <object>in sports competitions. Answer: <subject>
P451
unmarried partner
Fill in the blank: <subject>is the unmarried partner of ____. Answer: <object>
P54
member of sports team
@@
Fill in the blank: <subject>is a member of the sports team ____. Answer: <object>
P800
notable work
@@
Fill in the blank: <subject>’s notable work includes ____. Answer: <object>
P551
residence
Fill in the blank: <subject>’s residence is in ____. Answer: <object>
P131
located in the administrative territorial entity
@@
Fill in the blank: <subject>is located in the administrative territorial entity ____. Answer: <object>
P106
occupation
Fill in the blank: <subject>’s occupation is ____. Answer: <object>
P69
educated at
Fill in the blank: <subject>was educated at ____. Answer: <object>
P509
cause of death
Fill in the blank: <subject>’s cause of death was ____. Answer: <object>
P102
member of political party
@@
Fill in the blank: <subject>is a member of the political party ____. Answer: <object>
P19
place of birth
Fill in the blank: <subject>’s place of birth is ____. Answer: <object>
P115
home venue
Fill in the blank: <subject>’s home venue is ____. Answer: <object>
P1001
applies to jurisdiction
Fill in the blank: <object>applies to the jurisdiction of ____. Answer: <subject>
P840
narrative location
Fill in the blank: <subject>is set in the narrative location of ____. Answer: <object>
P108
employer
Fill in the blank: <subject>is employed by ____. Answer: <object>
P57
director
Fill in the blank: ____ is the director of <object>. Answer: <subject>
P2416
sports discipline competed in
@@
Fill in the blank: <subject>competes in the sports discipline of ____. Answer: <object>
P400
platform
Fill in the blank: <subject>is available on the platform ____. Answer: <object>
P1433
published in
Fill in the blank: <subject>was published in ____. Answer: <object>
P1056
product or material produced
@@
Fill in the blank: <subject>produces ____ as a product or material. Answer: <object>
P9071
character type
@@
Fill in the blank: <subject>is characterized as a ____ type. Answer: <object>
P4100
parliamentary group
@@
Fill in the blank: <subject>is a member of the parliamentary group ____. Answer: <object>
P937
work location
Fill in the blank: <subject>’s work location is ____. Answer: <object>
P1066
student of
Fill in the blank: <subject>is a student of ____. Answer: <object>
P1535
used by
Fill in the blank: <object>is used by ____. Answer: <subject>
P6
head of government
Fill in the blank: ____ is the head of government of <object>. Answer: <subject>
P2283
use
Fill in the blank: <subject>is used for ____. Answer: <object>
P812
academic major
Fill in the blank: <subject>’s academic major is ____. Answer: <object>
P1416
affiliation
@@
Fill in the blank: <subject>is affiliated with ____. Answer: <object>
P2522
victory
Fill in the blank: <subject>achieved a victory in ____. Answer: <object>
P607
conflict
Fill in the blank: <subject>was involved in the conflict ____. Answer: <object>
P749
parent organization
@@
Fill in the blank: <subject>is a part of the parent organization ____. Answer: <object>
P2283
uses
Fill in the blank: <subject>uses ____. Answer: <object>
P802
student
Fill in the blank: <subject>is a student at ____. Answer: <object>
P119
place of burial
Fill in the blank: <subject>’s place of burial is ____. Answer: <object>
P2842
place of marriage
@@
Fill in the blank: <subject>was married at ____. Answer: <object>
P286
head coach
Fill in the blank: ____ is the head coach of <object>. Answer: <subject>
P2541
operating area
Fill in the blank: <subject>’s operating area is ____. Answer: <object>
P1441
present in work
Fill in the blank: <subject>is present in the work ____. Answer: <object>
P2650
interested in
Fill in the blank: <subject>is interested in ____. Answer: <object>
P1027
conferred by
Fill in the blank: <object>is conferred by ____. Answer: <subject>
P3300
musical conductor
Fill in the blank: ____ is the musical conductor of <object>. Answer: <subject>
P2715
elected in
Fill in the blank: <subject>was elected in <object>. Answer: <object>
P2937
parliamentary term
@@
Fill in the blank: <subject>served during the parliamentary term ____. Answer: <object>
P1399
convicted of
Fill in the blank: <subject>was convicted of ____. Answer: <object>
P1686
for work
Fill in the blank: <subject>is used for the work ____. Answer: <object>
P1196
manner of death
Fill in the blank: <subject>’s manner of death was ____. Answer: <object>
P2632
place of detention
@@
Fill in the blank: <subject>was detained at ____. Answer: <object>
P991
successful candidate
Fill in the blank: <subject>was the successful candidate in ____. Answer: <object>
P2443
stage reached
Fill in the blank: <subject>reached the stage ____. Answer: <object>
P6872
has written for
Fill in the blank: <subject>has written for ____. Answer: <object>
@@
P5096
member of the crew of
Question: <subject>is a member of the crew of what? Answer: <object>.
P122
basic form of government
@@
Question: What is the basic form of government of <subject>? Answer: <object>.
P3448
stepparent
Question: <subject>is the stepparent of who? Answer: <object>.
P1479
has contributing factor
Question: <object>is a contributing factor to What? Answer: <subject>.
P61
discoverer or inventor
Question: Who is the discoverer or inventor of <object>? Answer: <subject>.
P3320
board member
@@
Question: <subject>serves as a board member of what? Answer: <object>.
P7779
member of military unit
@@
Question: <subject>is a member of the military unit of what? Answer: <object>.
P98
editor
Question: Who is is the editor of <object>? Answer: <subject>.
P1411
nominated for
Question: Who was nominated for <object>? Answer: <subject>.
P371
presenter
Question: Who is the presenter of <object>? Answer: <subject>.
P1365
replaces
Question: Whom did <subject>replace? Answer: <object>.
P488
chairperson
Question: Who serves as the chairperson of <object>? Answer: <subject>.
P27
country of citizenship
Question: What is <subject>’s country of citizenship? Answer: <object>.
P20
place of death
Question: Where is <subject>’s place of death? Answer: <object>.
P1344
participant in
Question: Who was a participant in <object>? Answer: <subject>.
P1366
replaced by
Question: What was <subject>replaced by? Answer: <object>.
P1412
languages spoken written or signed
@@
Question: What language does <subject>speaks? Answer: <object>.
P276
location
Question: Where is the location of <subject>? Answer: <object>.
P407
language of work or name
@@
Question: What is the language of <subject>’s work or name? Answer: <object>.
P39
position held

### v2

@@ 2 Related Work
Factual errors in an LLM’s output are also referred to as hallucinations (Xu et al., 2024) and occur when models generate content that is unfaithful to the provided input, established context, or factual correctness. Ji et al. (2023) categorises these into input-conflicting, context-conflicting, and fact-conflicting hallucinations, with the latter category being particularly problematic due to its contradiction with known facts.
While all scenarios in the PRobELM benchmark are technically fact-conflicting hallucinations (as they are not yet facts of the world the LLM has seen), the crucial point is that under existing world knowledge, non-factual scenarios are not all equally plausible. Instead, PRobELM probes a model’s ability to make educated guesses about plausible futures to gain insights into the capabilities of LLMs to infer within the limits of their knowledge beyond recalling facts it has already seen.
@@ 4.1 Experimental Setup
We first collect data points from Wikidata that align the timeframe of a given model (outlined in Table 1), following the steps detailed in section 3. For each most plausible scenario, we always add 10 less plausible alternatives, resulting in a total of 11 scenarios for models to rank. To ensure our evaluation is comprehensive yet with the practical constraints of time and computational resources, we sampled 126,000 scenarios (6,000 most plausible scenarios and 120,000 less plausible scenarios) from each timestamp, leading to four distinct timeframe-specific evaluation datasets. This sampling covered the five models under study, with OLMo and LLaMA models sharing the same timestamp. Details of the models and the corresponding timestamps are included in Appendix A.
Next, we follow the quality control measurements as detailed in subsection 3.3 to filter our data points. The final number of samples for each timestamp varies, as shown in Table 1.
@@ 5.1 Main Results on PRobELM
The results reveal that while model performance exceeded random chance, overall scores remain low, underscoring the complexity of the plausibility assessment task. For instance, considering a random baseline accuracy of 9.09% when ranking 11 scenarios, most models surpass this baseline across various sizes and prompts. Notably, both BLOOM models marginally outperform the random baseline by approximately 3 points, achieving 12.54% and 12.80% on the question-answering prompt. Specifically, the BLOOM-560M model underperform relative to the random baseline in the text completion prompt with 8.78% accuracy. This performance pattern suggests that language models face significant challenges in differentiating between scenarios of varying plausibility. It also highlights a distinct separation between models’ capabilities in factual accuracy and their ability to infer plausibility. These observations set the stage for our key findings:
@@
While variations in performance across prompts and metrics are observed, they typically correlate strongly. The most significant gap occur between OLMo-1B’s performance on the text completion prompt (22.11 acc.) and the question answering prompt (42.15 acc.), highlighting some variation in model proficiency across different types of tasks. Despite these variations, the assessment appears to hinge on a consistent underlying mechanism across prompt types. This is further evidenced by the Pearson correlation coefficients between the performances on various prompts, which were notably high: 0.8281 for statement versus text completion, 0.8672 for statement versus question answering, and 0.8976 for text completion versus question answering. Such strong correlations underscore that, despite the prompt-specific performance disparities, the models are evaluated based on a core aspect of their capability of ranking plausibility. Similarly, the Pearson correlation score between MRR and NDCG is 0.9982, indicating consistent plausibility measurements across metrics.
@@ 5.2 Comparison with Other Benchmarks
PRobELM shows distinct model performance compared to other reasoning benchmarks. Models that perform well on tasks emphasising factual accuracy do not always maintain their performance levels on plausibility assessments. In some cases, the performance rankings are inversely related. A notable example of this phenomenon is observed in the performance disparity between PRobELM and TruthfulQA. Specifically, Pythia-14M, which exhibits the best performance in TruthfulQA, positions near the bottom (9th place) of the PRobELM ranking. This variance underscores the distinct nature of plausibility as an evaluation criterion, demanding separate capabilities beyond those required for tasks focused on factual accuracy. The development of PRobELM thus contributes to the broader framework for evaluating language models by offering a different angle to assess their capability to understand and generate plausible content.
Figure 2: plausibility scores of models on the timeframe specific datasets. We illustrate one model per family. Vertical dotted lines represent the training timeframe of each model.
@@ 5.3 Temporal Effect of PRobELM
We further evaluate models against datasets from 2017, 2020, 2021, and 2023 timestamps to investigate how the recency of training data affects the plausibility performance. The outcomes of this analysis are presented in Figure 2 and detailed in Appendix C. Remarkably, the relative rankings of models prove relatively consistent across these varied timeframes, affirming the robustness of our evaluation approach. We observe that, from the 2020 to 2023 timestamps, models trained closer to or after the evaluation dates typically exhibit enhanced performance, underscoring the benefits of leveraging recent data.
@@ Appendix A Models Used in Our Experiments
(Touvron et al., 2023) LLaMA is recognised for its robust performance across a broad range of NLP tasks. Its training data, current as of August 2022, positions it uniquely for evaluation with the PRobELM benchmark, specifically designed to assess the model’s ability to infer plausibility based on its training up to that point. The closest corresponding timestamp we found is from December 2023 to January 2023.
@@ Appendix D More details on PRobELM
P5096 member of the crew of <subject>is a member of the crew of <object>.
P122 basic form of government The basic form of government of <subject>is <object>.
P3448 stepparent <subject>is the stepparent of <object>.
P1479 has contributing factor <object>is a contributing factor to <subject>.
P61 discoverer or inventor <subject>is the discoverer or inventor of <object>.
P3320 board member <subject>serves as a board member of <object>.
P7779 member of military unit <subject>is a member of the military unit <object>.
P98 editor <subject>is the editor of <object>.
P1411 nominated for <subject>was nominated for <object>.
P371 presenter <subject>is the presenter of <object>.
P1365 replaces <subject>replaces <object>.
P488 chairperson <subject>serves as the chairperson of <object>.
P27 country of citizenship <subject>’s country of citizenship is <object>.
P20 place of death <subject>’s place of death is <object>.
P1344 participant in <subject>was a participant in <object>.
P1366 replaced by <subject>was replaced by <object>.
@@
P1412 languages spoken written or signed <subject>speaks the following languages: <object>.
P276 location The location of <subject>is <object>.
P407 language of work or name The language of <subject>’s work or name is <object>.
P39 position held <subject>holds the position of <object>.
P1532 country for sport <subject>represents <object>in sports competitions.
P451 unmarried partner <subject>is the unmarried partner of <object>.
P54 member of sports team <subject>is a member of the sports team <object>.
P800 notable work <subject>’s notable work includes <object>.
P551 residence <subject>’s residence is in <object>.
P131 located in the administrative territorial entity <subject>is located in the administrative territorial entity <object>.
P106 occupation <subject>’s occupation is <object>.
P69 educated at <subject>was educated at <object>.
P509 cause of death <subject>’s cause of death was <object>.
P102 member of political party <subject>is a member of the political party <object>.
P19 place of birth <subject>’s place of birth is <object>.
P115 home venue <subject>’s home venue is <object>.
P1001 applies to jurisdiction <object>applies to the jurisdiction of <subject>.
P840 narrative location <subject>is set in the narrative location of <object>.
P108 employer <subject>is employed by <object>.
P57 director <subject>is the director of <object>.
P2416 sports discipline competed in <subject>competes in the sports discipline of <object>.
P400 platform <subject>is available on the platform <object>.
P1433 published in <subject>was published in <object>.
P1056 product or material produced <subject>produces <object>as a product or material.
P9071 character type <subject>is characterized as a <object>type.
P4100 parliamentary group <subject>is a member of the parliamentary group <object>.
P937 work location <subject>’s work location is <object>.
P1066 student of <subject>is a student of <object>.
P1535 used by <object>is used by <subject>.
P6 head of government <subject>is the head of government of <object>.
P2283 use <subject>is used for <object>.
P812 academic major <subject>’s academic major is <object>.
P1416 affiliation <subject>is affiliated with <object>.
P2522 victory <subject>achieved a victory in <object>.
P607 conflict <subject>was involved in the conflict <object>.
P749 parent organization <subject>is a part of the parent organization <object>.
P2283 uses <subject>uses <object>.
P802 student <subject>is a student at <object>.
P119 place of burial <subject>’s place of burial is <object>.
P2842 place of marriage <subject>was married at <object>.
P286 head coach <subject>is the head coach of <object>.
P2541 operating area <subject>’s operating area is <object>.
P1441 present in work <subject>is present in the work <object>.
P2650 interested in <subject>is interested in <object>.
P1027 conferred by <object>is conferred by <subject>.
P3300 musical conductor <subject>is the musical conductor of <object>.
P2715 elected in <subject>was elected in <object>.
P2937 parliamentary term <subject>served during the parliamentary term <object>.
P1399 convicted of <subject>was convicted of <object>.
P1686 for work <subject>is used for the work <object>.
P1196 manner of death <subject>’s manner of death was <object>.
P2632 place of detention <subject>was detained at <object>.
P991 successful candidate <subject>was the successful candidate in <object>.
P2443 stage reached <subject>reached the stage <object>.
P6872 has written for <subject>has written for <object>.
@@
P5096 member of the crew of Fill in the blank: <subject>is a member of the crew of ____. Answer: <object>
@@
P122 basic form of government Fill in the blank: The basic form of government of <subject>is ____. Answer: <object>
P3448 stepparent Fill in the blank: <subject>is the stepparent of ____. Answer: <object>
P1479 has contributing factor Fill in the blank: <object>is a contributing factor to ____. Answer: <subject>
P61 discoverer or inventor Fill in the blank: ____ is the discoverer or inventor of <object>. Answer: <subject>
@@
P3320 board member Fill in the blank: <subject>serves as a board member of ____. Answer: <object>
@@
P7779 member of military unit Fill in the blank: <subject>is a member of the military unit ____. Answer: <object>
P98 editor Fill in the blank: ____ is the editor of <object>. Answer: <subject>
P1411 nominated for Fill in the blank: ____ was nominated for <object>. Answer: <subject>
P371 presenter Fill in the blank: ____ is the presenter of <object>. Answer: <subject>
P1365 replaces Fill in the blank: <subject>replaces ____. Answer: <object>
P488 chairperson Fill in the blank: ____ serves as the chairperson of <object>. Answer: <subject>
P27 country of citizenship Fill in the blank: <subject>’s country of citizenship is ____. Answer: <object>
P20 place of death Fill in the blank: <subject>’s place of death is ____. Answer: <object>
P1344 participant in Fill in the blank: ____ was a participant in <object>. Answer: <subject>
P1366 replaced by Fill in the blank: <subject>was replaced by ____. Answer: <object>
@@
P1412 languages spoken written or signed Fill in the blank: <subject>speaks the following languages: ____. Answer: <object>
P276 location Fill in the blank: The location of <subject>is ____. Answer: <object>
@@
P407 language of work or name Fill in the blank: The language of <subject>’s work or name is ____. Answer: <object>
@@
P39 position held Fill in the blank: <subject>holds the position of ____. Answer: <object>
@@
P1532 country for sport Fill in the blank: ____ represents <object>in sports competitions. Answer: <subject>
P451 unmarried partner Fill in the blank: <subject>is the unmarried partner of ____. Answer: <object>
@@
P54 member of sports team Fill in the blank: <subject>is a member of the sports team ____. Answer: <object>
@@
P800 notable work Fill in the blank: <subject>’s notable work includes ____. Answer: <object>
P551 residence Fill in the blank: <subject>’s residence is in ____. Answer: <object>
@@
P131 located in the administrative territorial entity Fill in the blank: <subject>is located in the administrative territorial entity ____. Answer: <object>
P106 occupation Fill in the blank: <subject>’s occupation is ____. Answer: <object>
P69 educated at Fill in the blank: <subject>was educated at ____. Answer: <object>
P509 cause of death Fill in the blank: <subject>’s cause of death was ____. Answer: <object>
@@
P102 member of political party Fill in the blank: <subject>is a member of the political party ____. Answer: <object>
P19 place of birth Fill in the blank: <subject>’s place of birth is ____. Answer: <object>
P115 home venue Fill in the blank: <subject>’s home venue is ____. Answer: <object>
P1001 applies to jurisdiction Fill in the blank: <object>applies to the jurisdiction of ____. Answer: <subject>
P840 narrative location Fill in the blank: <subject>is set in the narrative location of ____. Answer: <object>
P108 employer Fill in the blank: <subject>is employed by ____. Answer: <object>
P57 director Fill in the blank: ____ is the director of <object>. Answer: <subject>
@@
P2416 sports discipline competed in Fill in the blank: <subject>competes in the sports discipline of ____. Answer: <object>
P400 platform Fill in the blank: <subject>is available on the platform ____. Answer: <object>
P1433 published in Fill in the blank: <subject>was published in ____. Answer: <object>
@@
P1056 product or material produced Fill in the blank: <subject>produces ____ as a product or material. Answer: <object>
@@
P9071 character type Fill in the blank: <subject>is characterized as a ____ type. Answer: <object>
@@
P4100 parliamentary group Fill in the blank: <subject>is a member of the parliamentary group ____. Answer: <object>
P937 work location Fill in the blank: <subject>’s work location is ____. Answer: <object>
P1066 student of Fill in the blank: <subject>is a student of ____. Answer: <object>
P1535 used by Fill in the blank: <object>is used by ____. Answer: <subject>
P6 head of government Fill in the blank: ____ is the head of government of <object>. Answer: <subject>
P2283 use Fill in the blank: <subject>is used for ____. Answer: <object>
P812 academic major Fill in the blank: <subject>’s academic major is ____. Answer: <object>
@@
P1416 affiliation Fill in the blank: <subject>is affiliated with ____. Answer: <object>
P2522 victory Fill in the blank: <subject>achieved a victory in ____. Answer: <object>
P607 conflict Fill in the blank: <subject>was involved in the conflict ____. Answer: <object>
@@
P749 parent organization Fill in the blank: <subject>is a part of the parent organization ____. Answer: <object>
P2283 uses Fill in the blank: <subject>uses ____. Answer: <object>
P802 student Fill in the blank: <subject>is a student at ____. Answer: <object>
P119 place of burial Fill in the blank: <subject>’s place of burial is ____. Answer: <object>
@@
P2842 place of marriage Fill in the blank: <subject>was married at ____. Answer: <object>
P286 head coach Fill in the blank: ____ is the head coach of <object>. Answer: <subject>
P2541 operating area Fill in the blank: <subject>’s operating area is ____. Answer: <object>
P1441 present in work Fill in the blank: <subject>is present in the work ____. Answer: <object>
P2650 interested in Fill in the blank: <subject>is interested in ____. Answer: <object>
P1027 conferred by Fill in the blank: <object>is conferred by ____. Answer: <subject>
P3300 musical conductor Fill in the blank: ____ is the musical conductor of <object>. Answer: <subject>
P2715 elected in Fill in the blank: <subject>was elected in <object>. Answer: <object>
@@
P2937 parliamentary term Fill in the blank: <subject>served during the parliamentary term ____. Answer: <object>
P1399 convicted of Fill in the blank: <subject>was convicted of ____. Answer: <object>
P1686 for work Fill in the blank: <subject>is used for the work ____. Answer: <object>
P1196 manner of death Fill in the blank: <subject>’s manner of death was ____. Answer: <object>
@@
P2632 place of detention Fill in the blank: <subject>was detained at ____. Answer: <object>
P991 successful candidate Fill in the blank: <subject>was the successful candidate in ____. Answer: <object>
P2443 stage reached Fill in the blank: <subject>reached the stage ____. Answer: <object>
P6872 has written for Fill in the blank: <subject>has written for ____. Answer: <object>
@@
P5096 member of the crew of Question: <subject>is a member of the crew of what? Answer: <object>.
@@
P122 basic form of government Question: What is the basic form of government of <subject>? Answer: <object>.
P3448 stepparent Question: <subject>is the stepparent of who? Answer: <object>.
P1479 has contributing factor Question: <object>is a contributing factor to What? Answer: <subject>.
P61 discoverer or inventor Question: Who is the discoverer or inventor of <object>? Answer: <subject>.
@@
P3320 board member Question: <subject>serves as a board member of what? Answer: <object>.
@@
P7779 member of military unit Question: <subject>is a member of the military unit of what? Answer: <object>.
P98 editor Question: Who is is the editor of <object>? Answer: <subject>.
P1411 nominated for Question: Who was nominated for <object>? Answer: <subject>.
P371 presenter Question: Who is the presenter of <object>? Answer: <subject>.
P1365 replaces Question: Whom did <subject>replace? Answer: <object>.
P488 chairperson Question: Who serves as the chairperson of <object>? Answer: <subject>.
P27 country of citizenship Question: What is <subject>’s country of citizenship? Answer: <object>.
P20 place of death Question: Where is <subject>’s place of death? Answer: <object>.
P1344 participant in Question: Who was a participant in <object>? Answer: <subject>.
P1366 replaced by Question: What was <subject>replaced by? Answer: <object>.
@@
P1412 languages spoken written or signed Question: What language does <subject>speaks? Answer: <object>.
P276 location Question: Where is the location of <subject>? Answer: <object>.
@@
P407 language of work or name Question: What is the language of <subject>’s work or name? Answer: <object>.